# Generalizing Goodstein’s theorem and Cichon’s independence proof

### v1

@@ Abstract
We generalize Goodstein’s theorem [2] and Cichon’s independence proof [1] to $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$ using results from [3]. The method is generalizable to stronger notation systems that provide unique terms for ordinals and enjoy Bachmann property.
@@ 1 Introduction
Assuming the familiar system $({\mathcal{E}},\cdot[\cdot])$ of fundamental sequences (limit approximations) for $\mathcal{E}$} and the definition of the slow-growing hierarchy $\mathrm{G}_{k}$} from [1], given by
${\mathrm{G}_{k}}(0):=0,\>{\mathrm{G}_{k}}(\alpha+1):={\mathrm{G}_{k}}(\alpha)+1,\mbox{ and }{\mathrm{G}_{k}}(\lambda):={\mathrm{G}_{k}}(\lambda[k])\mbox{ for }\lambda\in{\mathcal{E}}\cap\mathrm{Lim},$ (1)
as shown in [1] the restricted mapping
@@
is a bijection via transforming back and forth the bases $k$ and $\omega$, that is, writing a natural number $N$ in hereditary base-$k$ representation using 0, addition, and exponentiation to base $k$, replacing the base $k$ by $\omega$ we obtain ${\mathrm{G}_{k}^{-1}}(N)$, the unique preimage in $\nicefrac{{{\mathcal{E}}}}{{k}}$}. The image of $\nicefrac{{{\mathcal{E}}}}{{k}}$} under $\mathrm{G}_{k}$} is the Mostowski collapse of $\nicefrac{{{\mathcal{E}}}}{{k}}$} and equal to $\omega$, i.e., the restriction of $\mathrm{G}_{k}$} to $\nicefrac{{{\mathcal{E}}}}{{k}}$} is an order isomorphism.
@@
In this article we generalize this mechanism to the notation system $\operatorname{T}$} for ordinals below Takeuti ordinal, that is, to strength $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$, using fundamental sequences and machinery introduced in [3]. The method is generalizable to stronger notation systems that provide unique terms for ordinals and enjoy Bachmann property.
Adopting the preliminaries of [3], subsections 2.1, 2.3, and 2.5, which are summarized here in subsection 2.1 for reader’s convenience, let $\operatorname{T}$} be the notation system $\operatorname{T}^{\tau}$} for $\tau=1$ introduced there and let $\mathring{{\operatorname{T}}}$} be the subset of terms of countable cofinality. We called the system $({\mathring{{\operatorname{T}}}},\cdot[\cdot])$ of fundamental sequences with Bachmann property given in Definition 3.5 of [3] a Buchholz system. Bachmann property is a nesting property of fundamental sequences, namely that for terms $\alpha$ and $\beta$ such that $\alpha[\zeta]<\beta<\alpha$ and that are not regular cardinals, we have $\alpha[\zeta]\leq\beta[0]$, cf. Lemma 4.1 and Theorem 4.2 of [3].
For compatibility reasons with [1] we need to modify the definition of fundamental sequences in [3] so that for ordinals in the class of additive principal numbers, i.e. ordinals in the image of $\omega$-exponentiation $\xi\mapsto\omega^{\xi}$, say $\eta$, we redefine $(\eta\cdot\omega)[k]:=\eta\cdot k$, rather than $(\eta\cdot\omega)[k]:=\eta\cdot(k+1)$. [^1] Note that with this modification Bachmann property then only holds with appropriate restriction, which we will address later.
We will be able to use the approach by Cichon [1], transferring the definition of collapsing functions $\mathrm{G}_{k}$} as defined in (1) and predecessor functions $\operatorname{P}_{k}$}, defined as in [1] by
${\operatorname{P}_{k}}(0):=0,\>{\operatorname{P}_{k}}(\alpha+1):=\alpha,\mbox{ and }{\operatorname{P}_{k}}(\lambda):={\operatorname{P}_{k}}(\lambda[k])\mbox{ for }\lambda\in\mathrm{Lim},$ (2)
to the (modified) Buchholz system $({\mathring{{\operatorname{T}}}},\cdot[\cdot])$, more specifically to its countable initial segment. Note first of all that Cichon’s crucial lemma also holds in our context, with the same proof by induction on $\alpha$:
@@
These preparations allow us to quite canonically generalize Goodstein’s theorem [2] and Cichon’s independence proof in [1] to obtain independence of the theory $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$, the biggest of the “big five” theories in Simpson’s book on reverse mathematics. The argumentation, “Cichon’s trick”, is as follows: Given a starting base $k\in[2,\omega)$ and a natural number $N\in{\mathbb{N}}$, by Theorem 2.14 we obtain a unique $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$ such that
@@ 2.1 Preliminaries
We denote the class of limit ordinals as $\mathrm{Lim}$, the class of additive principal numbers as $\mathbb{P}$}, and the class of (non-zero) ordinals closed under $\omega$-exponentiation, also called $\varepsilon$-numbers, by $\mathbb{E}$}.
For terms $\alpha$ we use two abbreviations: first, writing $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$ means that $\eta\in{\mathbb{P}}$ and $\xi>0$ is minimal such that $\alpha=\xi+\eta$, second, $\alpha=_{\mathrm{\scriptscriptstyle{ANF}}}\xi_{1}+\ldots+\xi_{k}$ means that $\xi_{1},\ldots,\xi_{k}\in{\mathbb{P}}$ and $\xi_{1}\geq\ldots\geq\xi_{k}$. To indicate Cantor normal form representation of an ordinal $\alpha$, we write $\alpha=_{\mathrm{\scriptscriptstyle{CNF}}}\omega^{\xi_{1}}+\ldots+\omega^{\xi_{k}}$ if $\xi_{1}\geq\ldots\geq\xi_{k}$ and $k<\omega$. For the result of $l$-fold addition of the same term $\eta$ to a term $\xi$, where $l<\omega$, we will sometimes use the shorthand $\xi+\eta\cdot l$.
Terms in $\operatorname{T}$} are built up from 0, ordinal addition, and functions $\vartheta_{i}$ where $i<\omega$, where $\vartheta_{i}(0)={\Omega_{i}}$ and arguments of $\vartheta_{i}$ are restricted to terms below $\Omega_{i+2}$, setting $\Omega_{0}:=1$ and $\Omega_{i+1}:=\aleph_{i+1}$. As in subsection 2.1 of [3], slightly abusing notation we may consider notation systems ${\operatorname{T}}^{\Omega_{i+1}}$ to be systems relativized to the initial segment $\Omega_{i+1}$ of ordinals and built up over $\Omega_{i+1}=\vartheta_{i+1}(0),\Omega_{i+2}=\vartheta_{i+2}(0),\ldots$, i.e. without renaming the indices of $\vartheta$-functions. In this terminology ${\operatorname{T}}={\operatorname{T}}^{\Omega_{0}}$, and the sequence $({\operatorname{T}}^{{\Omega_{i}}})_{i<\omega}$ can be seen as an increasing sequence with larger and larger initial segments of ordinals serving as parameters.
The operation $\cdot^{\star_{i}}$ searches a $\operatorname{T}$}-term for its $\vartheta_{i}$-subterm of largest ordinal value, but under the restriction to treat $\vartheta_{j}$-subterms for $j<i$ as atomic. If such largest $\vartheta_{i}$-subterm does not exist, $\cdot^{\star_{i}}$ returns 0. The functions $\vartheta_{i}$ are natural extensions of the fixed point free Veblen functions and hence injective with images contained in the intervals ${\Omega_{i}},\Omega_{i+1})$, respectively for $i<\omega$. For $\alpha=\vartheta_{i}(\xi)$ and $\beta=\vartheta_{i}(\eta)$ (where $\xi$ and $\eta$ are any elements of the domain of $\vartheta_{i}$) we have
@@
As in [3] the arguments of $\vartheta_{i}$-terms $\vartheta_{i}(\xi)$ are often written in a form $\vartheta_{i}(\Delta+\eta)$, i.e. $\xi=\Delta+\eta$. This by convention always indicates that $\Omega_{i+1}\mid\Delta$ (possibly $\Delta=0$) and $\eta<\Omega_{i+1}$. Informally, we sometimes call the multiple $\Delta$ of $\Omega_{i+1}$ the fixed point level of $\vartheta_{i}(\xi)$, and the predicate $F_{i}(\Delta,\eta)$ means that $\eta=\sup_{\sigma<\eta}\vartheta_{i}(\Delta+\sigma)$, which according to Proposition 2.6 of [3] holds if and only if $\eta$ is of a form $\vartheta_{i}(\Gamma+\rho)$ where $\Gamma>\Delta$ and $\eta>\Delta^{\star_{i}}$.
Any $\xi\in{\operatorname{T}}\cap{\mathbb{P}}$ is of a form $\vartheta_{i}(\Delta+\eta)$ for unique $i<\omega$, $\Delta\in{\operatorname{T}}\cap\Omega_{i+2}$ with $\Omega_{i+1}\mid\Delta$, and $\eta<\Omega_{i+1}$. We will sometimes write $\xi\cdot\omega$ as a shorthand for the additive principal successor of $\xi$ in $\operatorname{T}$}, instead of $\vartheta_{i}(\xi)$ in case of $\Delta>0$ and $\vartheta_{i}(\eta+1)$ if $\Delta=0$.
For terms $\alpha$ of a form $\vartheta_{i}(\Delta+\eta)$ the $\Omega_{i}$}-localization, cf. subsection 2.3 of [3], is a sequence ${\Omega_{i}}=\alpha_{0},\ldots,\alpha_{m}=\alpha$ ($m\geq 0$), in which for $\alpha>{\Omega_{i}}$ the terms $\alpha_{i}=\vartheta_{i}(\Delta_{i}+\eta_{i})$, $i=1,\ldots,m$, provide a strictly increasing sequence of $\vartheta_{i}$-subterms of $\alpha$ that are not in the scope of any $\vartheta_{j}$-function with $j<i$ (i.e. that are not collapsed), such that the sequence of fixed point levels $\Delta_{1},\ldots,\Delta_{m}$ is strictly decreasing and of maximal length. Localization therefore approximates the ordinal $\alpha$ from below in terms of decreasing fixed point levels.
Restricting ordinal addition occuring in $\operatorname{T}$} to summation in additive normal form, the terms in $\operatorname{T}$} uniquely identify ordinal numbers below Takeuti ordinal. The term algebra developed in [3] and here therefore is algebra of ordinals.
For reader’s convenience we recall the following simple but crucial definition of a characteristic function regarding uncountable moduli and then provide the concrete system of fundamental systems which is a slightly modified version of Definition 3.5 of [3] that fits our context.
@@
We define a characteristic function $\chi^{\Omega_{i+1}}:{\operatorname{T}}^{\Omega_{i+1}}\to\{0,1\}$ where $i<\omega$ by recursion on the build-up of ${\operatorname{T}}^{\Omega_{i+1}}$:
@@
Note that this includes setting the support term for $\vartheta_{0}(0)=1$ to 0, and, more importantly, that $\underline{\alpha}$ need not be a subterm of $\alpha$, namely when $\eta$ is a successor ordinal. However, we may consider $\underline{\alpha}$ to enter the inductively generated set of terms (e.g. $\operatorname{T}$}) before $\alpha$ since it is either 0, a subterm of $\alpha$, or it requires one generative step less (addition of $1$) when $\eta$ is a successor.
@@
    For $\alpha=\vartheta_{i}(\Delta+\eta)>1$ where $i<\omega$, noting that $d\leq i$, the definition then proceeds as follows.
@@
      $\alpha[\zeta]:=\vartheta_{i}(\Delta+\eta[\zeta]).$
@@
        $\alpha[\zeta]:=\left\{\begin{array}[]{cl}\underline{\alpha}\cdot\zeta&\mbox{ if }\eta>0\mbox{ (and hence $d=0$)}\\ \zeta&\mbox{ otherwise.}\end{array}\right$.
@@
Bachmann property for the above notation system holds in the restricted sense that for terms $\alpha,\beta$ such that $\alpha[\zeta]<\beta<\alpha$ and that are not regular cardinals, we have $\alpha[\zeta]\leq\beta[1]$, where $\zeta<\aleph_{d}$ and $d=\mathrm{d}(\alpha)$. Note first that the base cases of the recursive definition of fundamental sequences of limit ordinals are clauses 3.2.1, covering additive principal successors and regular cardinals, and 3.2.2, covering successor fixed points. The modification we made therefore only affects limits of countable cofinality the fundamental sequence of which is recursively based on clause 3.2.1. Clause 3.2.2 is the only case where a fundamental sequence is defined in terms of another fundamental sequence of different cofinality. Thus, given $\alpha$, $\beta$ such that $\alpha[\zeta]<\beta<\alpha$ and that are not regular cardinals, we still have $\alpha[\zeta]\leq\beta[0]$, unless the fundamental sequence for $\beta$ is recursively based on clause 3.2.1, in which case $\alpha[\zeta]\leq\beta[1]$ still holds.
2.2 Quotients
For any $\beta\in{\operatorname{T}}$ it can be syntactically detected whether $\beta$ is of a form $\beta=\alpha[n]$ for some $\alpha\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$ and $n\in[2,\omega)$. Before we establish this explicitly, we use it in the following
Definition 2.4
  1. 1.
    $0\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$.
  2. 2.
  3. 3.
    If $\xi\in{\nicefrac{{{\operatorname{T}}}}{{k}}}\cap\Omega_{i+2}$ where $i<\omega$, then $\beta:=\vartheta_{i}(\xi)\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$, unless:
    1. (a)
      $\underline{\beta}\not\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$, or
    2. (b)
      $\beta$ is of a form $\alpha[n]$ where $\alpha\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$ and $n\geq k$.
[…]
Remark 2.5
It is easy to see that we have
[…]
By means of the following inversion lemma we will be able to recover $\beta$ from terms $\alpha\in{\operatorname{T}}$ of a form $\beta[n]$ where $n\in[2,\omega)$ and $\beta\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, as used in Definition 2.4. As in Definition 3.5 of [3] we need to formulate the inversion lemma more generally, so as to cover terms of uncountable cofinality. The inversion is formulated in a way that will cover the cases needed to define the quotients of Definition 2.4, rather than covering all possible cases.
Note that we rely on the modified version Definition 3.5 of [3] of fundamental sequences for $\operatorname{T}$}, redefining $\eta\cdot\omega[n]:=\eta\cdot n$ instead of $\eta\cdot(n+1)$ for $\eta\in{\mathbb{P}}$ as mentioned in the introduction.
@@ 2.2 Quotients
    $\zeta\in[2,\omega)$ and $\beta\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, or
@@
    $\zeta\in[{\Omega_{i}},\Omega_{i+1})\cap{\mathbb{P}}^{>1}$ for some $i<\omega$ and $\beta\in{\operatorname{T}}$ such that $\chi^{\Omega_{i+1}}(\beta)=1$
@@
    $\alpha=\zeta\in[{\Omega_{i}},\Omega_{i+1})\cap{\mathbb{P}}^{>1}$ and $\beta=\Omega_{i+1}=\vartheta_{i+1}(0)$.
@@
    $\alpha=\eta\cdot n$ for some $\eta\in{\operatorname{T}}\cap{\mathbb{P}}$ and $\zeta=n\in[2,\omega)$. Then $\beta:=\eta\cdot\omega$, and we have $\underline{\beta}=\eta$ and $\alpha=\beta[\zeta]$.
@@
    $\alpha=_{\mathrm{\scriptscriptstyle{ANF}}}\xi_{1}+\ldots+\xi_{k}+\rho$ for some $k\geq 1$ and $\rho=\eta[\zeta]$ according to the lemma’s conditions where $\eta\in(1,\xi_{k}]\cap{\mathbb{P}}$ so that $\rho<\xi_{k}$. Then setting $\beta:=\xi_{1}+\ldots+\xi_{k}+\eta$ we have $\alpha=\beta[\zeta]$.
@@
      $\rho$ is of a form $\rho=\eta[\zeta]$ according to the lemma’s conditions and $F_{j}(\Gamma,\eta)$ does not hold. Then setting $\Delta:=\Gamma$ and $\beta:=\vartheta_{j}(\Delta+\eta)$ we have $\alpha=\beta[\zeta]=\vartheta_{j}(\Delta+\eta[\zeta])$.

### v2

@@ Abstract
We generalize Goodstein’s theorem [2] and Cichon’s independence proof [1] to $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$ using results from Wilken [3]. The method is generalizable to stronger notation systems that provide or enable unique terms for ordinals and enjoy Bachmann property.  [^1] MSC: 03F40, 03D20, 03D60, 03F15.
@@ 1 Introduction
Defining the system $({\mathcal{E}},\cdot[\cdot])$ of fundamental sequences (limit approximations) for $\mathcal{E}$} by
$0[k]:=0,\;\;1[k]:=0,\;\;(\xi+\eta)[k]:=\xi+\eta[k],\;\;\omega^{\beta+1}[k]:=\omega^{\beta}\cdot(k+1),\;\mbox{ and }\;\omega^{\lambda}[k]:=\omega^{\lambda[k]},$
and the slow-growing hierarchy $\mathrm{G}_{k}$} for $k>0$ by [^2] Induction on $\alpha$ shows that our definitions are equivalent to Cichon’s in [1] for $k>0$, where Cichon’s $\alpha[k]$ is equal to our $\alpha[k-1]$.
${\mathrm{G}_{k}}(0):=0,\>{\mathrm{G}_{k}}(\alpha+1):={\mathrm{G}_{k}}(\alpha)+1,\mbox{ and }{\mathrm{G}_{k}}(\lambda):={\mathrm{G}_{k}}(\lambda[k-1])\mbox{ for }\lambda\in{\mathcal{E}}\cap\mathrm{Lim},$ (1)
where $\mathrm{Lim}$ denotes the class of limit ordinals, as shown in [1] the restricted mapping
@@
is a bijection via transforming back and forth the bases $k$ and $\omega$, that is, writing a natural number $N$ in hereditary base-$k$ representation using 0, addition, and exponentiation to base $k$, replacing the base $k$ by $\omega$ we obtain ${\mathrm{G}_{k}^{-1}}(N)$, the unique preimage of $N$ in $\nicefrac{{{\mathcal{E}}}}{{k}}$}. The image of $\nicefrac{{{\mathcal{E}}}}{{k}}$} under $\mathrm{G}_{k}$} is the Mostowski collapse of $\nicefrac{{{\mathcal{E}}}}{{k}}$} and equal to $\omega$, i.e., the restriction of $\mathrm{G}_{k}$} to $\nicefrac{{{\mathcal{E}}}}{{k}}$} is an order isomorphism.
@@
In this article we generalize this mechanism to the notation system $\operatorname{T}$} for ordinals below Takeuti ordinal, that is, to the strength of the fragment $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$ of second order number theory, which is the strongest of the “big five” theories in Simpson’s book on reverse mathematics, using fundamental sequences and machinery introduced in [3]. The method is generalizable to stronger notation systems that provide or enable unique terms for ordinals and enjoy Bachmann property.
Adopting the preliminaries of [3], subsections 2.1, 2.3, and 2.5, which are summarized here in subsection 2.1 for reader’s convenience, let $\operatorname{T}$} be the notation system $\operatorname{T}^{\tau}$} for $\tau=1$ introduced there and let $\mathring{{\operatorname{T}}}$} be the subset of terms of countable cofinality. The system $({\mathring{{\operatorname{T}}}},\cdot[\cdot])$ of fundamental sequences with Bachmann property given in Definition 3.5 of [3] is called a Buchholz system. Bachmann property is a nesting property of fundamental sequences, namely that for terms $\alpha$ and $\beta$ such that $\alpha[\zeta]<\beta<\alpha$ where $\alpha$ is not a successor-multiple of a regular cardinal $\Omega_{i+1}$, we have $\alpha[\zeta]\leq\beta[0]$, see Theorem 2.6.
We will be able to use the approach by Cichon [1], transferring the definition of collapsing functions $\mathrm{G}_{k}$} as defined in (1) and predecessor functions $\operatorname{P}_{k}$} for $k>0$, defined equivalently as in [1] by
${\operatorname{P}_{k}}(0):=0,\>{\operatorname{P}_{k}}(\alpha+1):=\alpha,\mbox{ and }{\operatorname{P}_{k}}(\lambda):={\operatorname{P}_{k}}(\lambda[k-1])\mbox{ for }\lambda\in\mathrm{Lim},$ (2)
to the Buchholz system $({\mathring{{\operatorname{T}}}},\cdot[\cdot])$, more specifically to its countable initial segment. Note first of all that Cichon’s crucial lemma also holds in our context, with the same proof by induction on $\alpha$:
@@
These preparations allow us to quite canonically generalize Goodstein’s theorem [2] and Cichon’s independence proof in [1] to obtain independence of the theory $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$. The argumentation, “Cichon’s trick”, is as follows: Given a starting base $k\in[2,\omega)$ and a natural number $N\in{\mathbb{N}}$, by Theorem 2.21 we obtain a unique $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$ such that
@@ 2.1 Preliminaries
Besides denoting the class of limit ordinals as $\mathrm{Lim}$, we denote the class of additive principal numbers as $\mathbb{P}$}, ${\mathbb{P}}^{>1}:={\mathbb{P}}\setminus\{1\}$, and the class of (non-zero) ordinals closed under $\omega$-exponentiation, also called $\varepsilon$-numbers, by $\mathbb{E}$}.
For terms $\alpha$ we use the following abbreviations. Writing $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$ means that $\eta\in{\mathbb{P}}$ and $\xi\geq\eta$ is minimal such that $\alpha=\xi+\eta$, $\alpha=_{\mathrm{\scriptscriptstyle{ANF}}}\xi_{1}+\ldots+\xi_{k}$ means that $\xi_{1},\ldots,\xi_{k}\in{\mathbb{P}}$ and $\xi_{1}\geq\ldots\geq\xi_{k}$, and to indicate Cantor normal form representation of an ordinal $\alpha$, we write $\alpha=_{\mathrm{\scriptscriptstyle{CNF}}}\omega^{\zeta_{1}}+\ldots+\omega^{\zeta_{k}}$ where $\zeta_{1}\geq\ldots\geq\zeta_{k}$. For the result of $l$-fold addition of the same term $\eta$ to a term $\xi$, where $l<\omega$, we will sometimes use the shorthand $\xi+\eta\cdot l$. We further define ${\mathrm{end}}(0):=0$ and ${\mathrm{end}}(\alpha):=\xi_{k}$ if $\alpha=_{\mathrm{\scriptscriptstyle{ANF}}}\xi_{1}+\ldots+\xi_{k}$ where $k\geq 1$.
Setting $\Omega_{0}:=\omega$ and $\Omega_{i+1}:=\aleph_{i+1}$, terms in $\operatorname{T}$} are built up from 0, 11, ordinal addition, and functions $\vartheta_{i}$ for each $i<\omega$, where $\vartheta_{i}(0)={\Omega_{i}}$ and arguments of $\vartheta_{i}$ are restricted to terms below $\Omega_{i+2}$. The functions $\vartheta_{i}$ are natural extensions of the fixed point-free Veblen functions and hence injective, with images contained in the intervals ${\Omega_{i}},\Omega_{i+1})$, respectively for $i<\omega$. Compared to the setup in [3], defining $\Omega_{0}:=\omega$ instead of $\Omega_{0}=1$ and consequently $\vartheta_{0}(i):=\omega^{i+1}$ instead of $\vartheta_{0}(i)=\omega^{i}$ for $i<\omega$ is the only modification, which is needed for technical smoothness.
As in subsection 2.1 of [3], slightly abusing notation we may consider notation systems ${\operatorname{T}}^{\Omega_{i+1}}$ to be systems relativized to the initial segment $\Omega_{i+1}$ of ordinals and built up over $\Omega_{i+1}=\vartheta_{i+1}(0),\Omega_{i+2}=\vartheta_{i+2}(0),\ldots$, i.e. without renaming the indices of $\vartheta$-functions. In this terminology ${\operatorname{T}}={\operatorname{T}}^{\Omega_{0}}$, and the sequence $({\operatorname{T}}^{{\Omega_{i}}})_{i<\omega}$ can be seen as an increasing sequence with larger and larger initial segments of ordinals serving as parameters.
The operation $\cdot^{\star_{i}}$ searches a $\operatorname{T}$}-term for its $\vartheta_{i}$-subterm of largest ordinal value, but under the restriction to treat $\vartheta_{j}$-subterms for $j<i$ as atomic. If such largest $\vartheta_{i}$-subterm does not exist, $\cdot^{\star_{i}}$ returns 0. For $\alpha=\vartheta_{i}(\xi)$ and $\beta=\vartheta_{i}(\eta)$ (where $\xi$ and $\eta$ are any elements of the domain of $\vartheta_{i}$) we have
@@
As in [3] the arguments of $\vartheta_{i}$-terms $\vartheta_{i}(\xi)$ are often written in a form $\vartheta_{i}(\Delta+\eta)$, i.e. $\xi=\Delta+\eta$. This by convention always indicates that $\Omega_{i+1}\mid\Delta$ (possibly $\Delta=0$) and $\eta<\Omega_{i+1}$. Informally, we sometimes call the multiple $\Delta$ of $\Omega_{i+1}$ the fixed point level of $\vartheta_{i}(\xi)$, and the predicate $F_{i}(\Delta,\eta)$ means that $\eta=\sup^{+}_{\sigma<\eta}\vartheta_{i}(\Delta+\sigma)$, which according to Proposition 2.6 of [3] holds if and only if $\eta$ is of a form $\vartheta_{i}(\Gamma+\rho)$ where $\Gamma>\Delta$ and $\eta>\Delta^{\star_{i}}$.
Any $\alpha\in{\operatorname{T}}\cap{\mathbb{P}}^{>1}$ is of a form $\vartheta_{i}(\Delta+\eta)$ for unique $i<\omega$, $\Delta\in{\operatorname{T}}\cap\Omega_{i+2}$ with $\Omega_{i+1}\mid\Delta$, and $\eta<\Omega_{i+1}$. Let $\operatorname{lv}(\alpha):=i$ denote the level of $\alpha$ and for $j\leq i$ let $\operatorname{Par}_{j}(\alpha)$ denote the set of $\vartheta$-subterms of level strictly less than $j$ (so that $\operatorname{Par}_{0}(\alpha)=\emptyset$). We will sometimes write $\alpha\cdot\omega$ as a shorthand for the additive principal successor of $\alpha=\vartheta_{i}(\Delta+\eta)$ in $\operatorname{T}$}, instead of $\vartheta_{i}(\alpha)$ in case of $\Delta>0$ and $\vartheta_{i}(\eta+1)$ if $\Delta=0$.
For terms $\alpha$ of a form $\vartheta_{i}(\Delta+\eta)$ the $\Omega_{i}$}-localization, cf. subsection 2.3 of [3], is a sequence ${\Omega_{i}}=\alpha_{0},\ldots,\alpha_{m}=\alpha$ ($m\geq 0$), in which for $\alpha>{\Omega_{i}}$ the terms $\alpha_{i}=\vartheta_{i}(\Delta_{i}+\eta_{i})$, $i=1,\ldots,m$, provide a strictly increasing sequence of $\vartheta_{i}$-subterms of $\alpha$ that are not in the scope of any $\vartheta_{j}$-function with $j<i$ (i.e. that are not collapsed), such that the sequence of fixed point levels $\Delta_{1},\ldots,\Delta_{m}$ is strictly decreasing and of maximal length, and where the $\eta_{i}$ are maximal. Localization therefore approximates the ordinal $\alpha$ from below in terms of decreasing fixed point levels.
Restricting ordinal addition occuring in $\operatorname{T}$} to summation in additive normal form, the terms in $\operatorname{T}$} uniquely identify ordinal numbers below Takeuti ordinal. The term algebra developed in [3] and here therefore can be understood as algebra of ordinals.
We introduce a simple but crucial term measure for inductive definitions along ordinals in $\operatorname{T}$}.
Definition 2.1
We define the height measure ${\operatorname{ht}_{i}}:{\operatorname{T}}\to\omega$ where $i<\omega$ as follows.
[…]
We may extend the definition of $\operatorname{ht}_{i}$} also to the set ${\operatorname{T}}^{{\Omega_{i}}}$ which instead of terms $\vartheta_{j}$ for $j<i$ contains opaque parameters for all ordinals below $\Omega_{i}$}.
Lemma 2.2
For $\alpha<{\operatorname{T}}^{{\Omega_{i}}}\cap\Omega_{i+1}$
${\operatorname{ht}_{i}}(\alpha)=\min\{n\>|\>\alpha<\vartheta_{i}(\ldots(\vartheta_{i+n}(0))\ldots)\}.$
$\operatorname{ht}_{i}$} is weakly increasing on ${\operatorname{T}}^{{\Omega_{i}}}\cap\Omega_{i+1}$.
For reader’s convenience we recall from [3] the following crucial definition of a characteristic function regarding uncountable moduli and then provide the concrete system of fundamental sequences adapted from Definition 3.5 of [3] to fit with our modified definition of $\vartheta_{0}(i)$ for $i<\omega$.
@@
We define characteristic functions $\chi^{\Omega_{i+1}}:{\operatorname{T}}^{\Omega_{i+1}}\to\{0,1\}$ for $i<\omega$ by recursion on the build-up of ${\operatorname{T}}^{\Omega_{i+1}}$:
@@
    For $\alpha=\vartheta_{i}(\Delta+\eta)$ where $i<\omega$, noting that $d\leq i$, the definition then proceeds as follows.
@@
      $\alpha[\zeta]:=\left\{\begin{array}[]{cl}\vartheta_{0}(\zeta)&\mbox{ if }\alpha=\vartheta_{0}(\vartheta_{0}(0))=\omega^{\omega}\\[5.69054pt] \vartheta_{i}(\Delta+\eta[\zeta])&\mbox{ otherwise.}\end{array}\right$.
@@
        $\alpha[\zeta]:=\left\{\begin{array}[]{cl}\underline{\alpha}\cdot(\zeta+1)&\mbox{ if }d=0,\\ \zeta&\mbox{ if }d>0.\end{array}\right$.
@@
The following theorem states that terms in $\operatorname{T}$} satisfy Bachmann property, which is crucial for our approach toward generalizing Goodstein’s theorem.
Theorem 2.6 (Lemma 4.1 and Theorem 4.2 of [3])
For $\alpha,\beta\in{\operatorname{T}}$, where $\alpha$ is not a successor-multiple of a regular cardinal, and for which $\alpha[\zeta]<\beta<\alpha$ holds for some $\zeta<\aleph_{\mathrm{d}(\alpha)}$, we have
$\alpha[\zeta]\leq\beta[0].$
The following lemma states a basic observation about fundamental sequences with Bachmann property. Note first that whenever $\alpha[\xi]=\alpha[\zeta]$ where $\alpha\in{\operatorname{T}}$ and $\xi,\zeta<\aleph_{\mathrm{d}(\alpha)}$, then $\xi=\zeta$ since fundamental sequences are strictly increasing. Note further that for any $\alpha\in{\operatorname{T}}\cap{\mathbb{P}}$ we have $\alpha\cdot\omega[0]=\alpha$.
Corollary 2.7
Let $\alpha,\beta\in{\operatorname{T}}$ such that $\beta<\alpha$ and $\alpha$ is not a successor-multiple of any $\Omega_{j+1}$. Let $\zeta<\aleph_{\mathrm{d}(\beta)}$, $\xi<\aleph_{\mathrm{d}(\alpha)}$, and suppose that
$\alpha[\xi]=\beta[\zeta].$
Then it follows that $\zeta=0$.
Proof. This is an immediate consequence of Bachmann property, Theorem 2.6. $\Box$
By means of the following inversion lemma we will be able to recover $\beta$ from terms $\alpha\in{\operatorname{T}}$ of a form $\beta[n]$ where $n\in[1,\omega)$ and $\beta\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, as used in Definition 2.12. As in Definition 3.5 of [3] we need to formulate the inversion lemma more generally, so as to cover terms of uncountable cofinality. The inversion is formulated in a way that will cover the cases needed to define the quotients of Definition 2.12, rather than covering all possible cases.
@@
    $\zeta\in[1,\omega)$ and $\beta\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, or
@@
    $\zeta\in[{\Omega_{i}},\Omega_{i+1})\cap{\mathbb{P}}$ for some $i<\omega$ and $\beta\in{\operatorname{T}}$ such that $\chi^{\Omega_{i+1}}(\beta)=1$
@@
    $\alpha=\zeta\in[{\Omega_{i}},\Omega_{i+1})\cap{\mathbb{P}}$ and $\beta=\Omega_{i+1}=\vartheta_{i+1}(0)$.
@@
    $\alpha=\eta\cdot(n+1)$ for some $\eta\in{\operatorname{T}}\cap{\mathbb{P}}$ and $\zeta=n\in[1,\omega)$. Then $\beta:=\eta\cdot\omega$, and we have $\underline{\beta}=\eta$ and $\alpha=\beta[\zeta]$.
@@
    $\alpha=\xi_{1}+\ldots+\xi_{k}+\rho$ where $\xi_{1},\ldots,\xi_{k}\in{\mathbb{P}}$ is weakly decreasing, $k\geq 1$, and $\rho=\eta[\zeta]$ according to the lemma’s conditions with $\eta\in(1,\xi_{k}]\cap{\mathbb{P}}$ so that $\rho<\xi_{k}$. Then setting $\beta:=\xi_{1}+\ldots+\xi_{k}+\eta$ we have $\alpha=\beta[\zeta]$.
@@
      $\rho$ is of a form $\rho=\eta[\zeta]$ according to the lemma’s conditions where $\eta<\Omega_{j+1}$ and $F_{j}(\Gamma,\eta)$ does not hold. Then setting $\Delta:=\Gamma$ and $\beta:=\vartheta_{j}(\Delta+\eta)$ we have $\alpha=\beta[\zeta]=\vartheta_{j}(\Delta+\eta[\zeta])$.